OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x2 − 4x − 60 −10, 6 −14, 5 −6, 10 −5, 14

OpenStudy (anonymous):

i guess this one factors try \[(x-10)(x+6)=0\] then set \[x-10=0\] and \[x+6=0\] and solve for \(x\)